# GUing: A Mobile GUI Search Engine using a Vision-Language Model

### v1

@@ 2.3.1. Object Detection Model
We use DETR (DEtection TRansformer) (Carion et al., 2020) as our screenshot detector. DETR is an end-to-end object detector, composed of a CNN (convolutional neural network) backbone and a encoder-decoder transformer (Vaswani et al., 2017). Atop the transformer decoder, dual output heads are appended: a linear layer dedicated to categorize class labels, and a multi-layer perceptron (MLP) tasked with the generation of bounding boxes for the object location. The model uses so-called object queries to detect objects in an image. Each object query is designed to search specifically for one particular object of interest in the given image. In our work, the number of queries is set to 1, as we are interested in the largest screenshot within a surrounded screenshot.
2.3.2. Training Data Creation
To fine tune the detection model for screenshot cropping, we created a dataset. From the dataset created in Section 2.2.2, we collected all the surrounded screenshots and labeled the screenshot areas with a bounding box. The annotation was performed with Prodigy (Explosion, 2024). For each image, an author drew one bounding boxes that cover the screenshot, the results were reviewed by another author. The majority of the surrounded screenshots feature only a single screenshot. In instances where a single image contains multiple screenshots, annotation was selectively applied only to the most prominent (largest) screenshot. This resulted in a dataset comprising 1,768 annotated images, each including a delineated bounding box that specifies the screenshot area.
@@ 2.3.3. Detection Performance
Intersection over Union (IoU) is a standard metric in object detection tasks. It evaluates the extent of overlap between the predicted bounding box and the ground truth bounding box. If $IoU=0$, it means that there is no overlap between the boxes, while $IoU=1$ means that the overlap is perfect.
@@
In this study, we followed the evaluation protocol established by the Common Objects in Context (COCO) dataset (Lin et al., 2014), which is a common benchmark in the field of object detection. As described in the protocol, we quantitatively evaluated the performance of our object detection by calculating precision and recall metrics at various IoU thresholds. Specifically, we computed the precision values at three distinct IoU levels: 0.50, 0.75, and an average over the range from 0.50 to 0.95. Similarly, recall was determined over the IoU range of 0.50 to 0.95.
We applied 80% of the data for training and the remaining 20% for evaluation. The detector was trained on the training set using mini-batch gradient descent, with a batch size of 16 and AdamW optimizer with an initial learning rate of $1e^{-5}$ and the training epochs is 10. Table 2 shows the results of ten-fold cross validation. The IoU at 0.75 exhibits a value of 0.971, indicating a high level of accuracy.
Table 2. Screen localization accuracy
@@ 2.3.4. Applying the Detector
We trained a DETR model with all of the 1,768 images, and applied this model to the 191,993 images that were categorized as surrounded screenshot in Section 2.2.4. DETR predicted a bounding box that localize the screenshot for each of the images. We then cropped these images according to the bounding box. The screenshots cropped from the surrounded screenshot as well as the app introduction images which were previously classified as screenshots represent the GPSRepo dataset.
@@ 2.4. Caption Recognition
The majority of the images, classified as surrounded screenshots, include captions that provide a succinct descriptions of the screenshots. Typically, these captions are positioned either above or below the image. In order to accurately extract this textual information from the images, we employed Optical Character Recognition (OCR) technology. This technology is adept at detecting and recognizing text displayed within image files.
@@ 2.4.1. Applying OCR
In this study, we employed PaddleOCR (Li et al., 2022b), which is an industrially robust OCR model renowned for its accuracy and swift performance. Given that the textual content within the images under analysis is markedly legible, exhibiting a clarity not typically associated with hand-written text, the application of PaddleOCR is deemed adequately proficient for this task. PaddleOCR processes an input image and outputs a collection of identified elements. Each element is comprised of a bounding box delineating the text area, the recognized text within, and an associated confidence level reflecting the model’s recognition certainty.
We applied PaddleOCR on all of the 191,993 surrounded screenshots. For each surrounded screenshot, text bounding boxes that exhibit spatial overlap with the screenshot’s bounding box, as delineated in Section 2.3.4, were excluded. The remaining text bounding boxes were subsequently aggregated to form a coherent caption for the corresponding image.
@@ 2.4.2. Post-processing
Some of the surrounded screenshots may not contain captions. These images were removed from our analysis. The captions extracted are not exclusively in English; they encompass a variety of languages including French, German, and Arabic, among others. As we focus on English search in this work, captions in languages other than English were excluded from the analysis. To facilitate the identification of the language to which each caption pertains, we employed Lingua (Stahl, 2022), an effective language detection tool. Furthermore, to correct any spelling error within the captions due to potential OCR errors, we utilized Autocorrect (Sondej, 2020), a Python-based spelling corrector.
It is noteworthy that in some instances, multiple surrounded screenshots of an app may feature identical captions. This repetition suggests that the caption is either overly generic or simply a logo of the app, and thus may not adequately convey the screenshot content. To address this redundancy, we conducted a filtering process whereby, for each app, only the first surrounded screenshot from the set containing identical captions was retained. We observed that the first surrounded screenshot displayed by an app on Google Play typically showcases the most representative app feature. After the post-processing stage, we obtained 135,357 screenshot-caption pairs, collectively referred to as the GPSCap dataset.
@@ 3. GUI Search Engine
We introduce GUing, an advanced search engine based on vision-language model that leverages textual queries to retrieve relevant screenshots from a large GUI repository which consists of GPSRepo, Rico (Deka et al., 2017) and ReDraw (Moran et al., 2020). The architecture of GUing is illustrated in Figure 4.
Figure 4. Overview of our GUI search engine (GUing) for text-to-GUI retrieval.
@@
Overview of our GUI search engine (GUing) for text-to-GUI retrieval.
GUing is developed based on the UIClip model. This search engine utilizes the image encoder and text encoder modules of UIClip to embed screenshots and textual queries within a unified latent space. Consequently, it becomes possible to facilitate an efficient search process for the screenshots by calculating the cosine similarity between the text query embedding and screenshot embedding. In the following, we elaborate on the details of our GUI search engine.
3.1. Constructing UIClip Model
UIClip, our vision-language model, serves as the fundamental component of our GUI search engine, integrating the image and text modalities. Building upon the foundation of CLIP (Radford et al., 2021), UIClip offers enhanced capabilities in multimodal representation learning within GUI domain. CLIP (Contrastive Language–Image Pre-training) model based on vision language contrastive learning on large-scale image-caption data is a significant foundation model in multimodal representation learning. In this work, we construct a GUI-specific CLIP model by training a vision-language model on a large-scale dataset of screenshot-caption pairs.
@@ 3.1.1. Training Data and Pre-processing
For the training of UIClip, we combined three datasets to learn the visual representation of mobile screenshots:
  • Screen2Words dataset (Wang et al., 2021) is a subset of Rico. It contains 112,085 language summarization across 22,417 unique screenshots. The screenshot summaries are written by professional annotators. For each screenshot, five summaries are created by five different annotators.
  • Clarity dataset (Moran et al., 2022) consists of 45,998 descriptions for 10,204 screenshots from popular Android apps. Crowd workers summarize each screenshot in different granularity, with one high-level caption to up to four low-level detailed descriptions.
  • Google Play Screenshot Caption (GPSCap) dataset. Here the textual description is created by the respective developers and app vendors. In the Section 2, we created a dataset with 135k screenshot-caption pairs by mining Google Play. The screenshots are cropped from app introduction images, while the corresponding captions describe briefly the screenshots.
The captions were firstly tokenised with the CLIP tokeniser (OpenAI, 2021a). Captions longer than 77 tokens were truncated to 77, captions shorter were padded to 77. In terms of image processing for the screenshots, the resolution was resized to $224\times 224$. Furthermore, the pixel value of the images was proportionately rescaled from a range of 0-255 to 0-1. This new value was subsequently normalized using the parameters (means and standard deviations) featured in the official CLIP implementation documentation (OpenAI, 2021a).
@@ 3.1.2. Model Architecture
UIClip consists of two components, an image encoder and a text encoder. The image encoder, which adopts a Vision Transformer (ViT) model (Dosovitskiy et al., 2021), ingests image data and generates its corresponding image embedding. The text encoder is a transformer encoder (Vaswani et al., 2017) which accepts text as input to produce a associated text embedding. Employing contrastive learning techniques, the image and text embeddings are mapped into a multi-modal embedding space. In this shared space, images and texts presenting semantic similarity are mapped close to each other.
Contrastive learning on image-text pairs constructs a bridge between visual perception and natural language. Given a batch of N (image, text) pairs, the training objective is to identify the authentic pairings from $N\times N$ potential (image, text) combinations within the batch (Radford et al., 2021). To achieve this objective, the model formulates a multi-modal embedding space, concurrently training an image encoder and text encoder to maximize the cosine similarity of the image and text embeddings derived from the N authentic pairs in the batch. Simultaneously, it minimizes the cosine similarity of the embeddings from the remaining $N^{2}-N$ incorrect pairings.
Figure 5. Overview of contrastive learning to train our vision-language model UIClip (adapted from (Radford et al., 2021))
@@
Overview of contrastive learning to train our vision-language model UIClip (adapted from (Radford et al., 2021)).
@@ 3.1.3. Training Details
In order to adapt the CLIP model to UIClip, rather than training the CLIP model from scratch, where the weights of image encoder and text encoder are randomly initialized, our training is based on the checkpoint of ”openai/clip-vit-base-patch32” (OpenAI, 2021b). This checkpoint has been pre-trained on a 400 million image-text pairs. This method allows us to harness the benefits inherent to pre-trained models and, as a result, substantially diminishes the time required for training.
The model was trained on all of the collected screenshot-caption pairs for 5 epochs. We applied mini-batch gradient descent with a batch size of 128 and AdamW optimizer with an initial learning rate of $5e^{-5}$.
@@ 3.2.1. Processing the GUI Repository
The GUI repository is composed of 383k screenshots from GPSRepo (303k), Rico (66k) and ReDraw (14k). These screenshots are on image format, thus cannot be used directly for querying. Therefore, we embeded all the screenshots in the GUI repository and save them in a vector database.
The screenshots are firstly resized, rescaled and normalized with the same parameters as those in Section 3.1.1. These processed screenshots subsequently serve as the input for the image encoder of UIClip. The image encoder accepts an image as an input, subsequently outputting a 512-dimensional embedding. The entire GUI repository, comprising 383k screenshots, is processed via the image encoder, consequently generating 383k 512-dimensional embeddings. The mapping between the screenshot id and the resulted embeddings are saved in a vector database for further retrieval.
@@ 3.2.2. Querying on the GUI Repository
In the querying phase, the textual query is encoded by UIClip’s text encoder, resulting in a 512-dimensional embedding. Subsequently, this query embedding is utilized to identify the top-k most similar screenshots embeddings from the vector database. In our research, we employ cosine similarity as the measure of similarity. Consequently, the query results consist of the top-k screenshots that bear the most resemblance.
@@ 3.2.3. Accelerating the Searching
Given the large scale of the GUI repository, employing a brute force approach to compare the query embedding with all screenshot embeddings is inefficient. To tackle this problem, we applied approximate nearest neighbor search (Arya et al., 1998) to accelerate the querying as described following, as detailed below.
We established n Voronoi cells within the embedding space, with each screenshot embedding falling into one of these designated cells. During a search, the query embedding is initially compared with the centroid embeddings of each cell. This method results in a considerable reduction of necessary comparisons when contrasted with the total number of screenshot embeddings. The k cells with centroids closest to the query embedding are identified, whereupon the screenshot embeddings within these cells are thereafter compared with the query embedding in order to pinpoint similar screenshot embeddings. Our implementation is based on Faiss (Douze et al., 2024), a library for efficient similarity search. We set the number of cells n to 3000, and the number of cells that are visited to perform a search k to 1000.
@@ 4. Evaluation Design
We present the design of the empirical evaluation for our GUI search engine. Specifically, we investigate the following research questions:
RQ1: How does GUing perform in text-to-GUI retrieval tasks compared to state of the art approaches? As the CLIP model (Radford et al., 2021) was trained on 400 millions of image-caption pairs from internet, it can also be applied on text-to-GUI retrieval task. An alternative method involves leveraging solely the textual information - the text displayed on the GUI images, computing its semantic similarity with the input textual query, and subsequently ranking this similarity for effective GUI retrieval. Thus, the question is whether the performance our proposed UIClip model surpasses the CLIP model and text-only approaches.
RQ2: How relevant are the retrieved screenshots for queries representing app features? As a search engine, GUing processes textual queries as input and searches for relevant images within the GUI repository. It is thus important to evaluate how effectiveness GUing in delivering useful results that actually align with user expectations.
To answer these research questions, we designed two experiments: One (Expt1) is based on benchmarking three screenshot-caption datasets, and the other (Expt2) is based on a manual evaluation of the search engines.
4.1. Expt1: Evaluation of Models Performance on Three Datasets
In the first experiment, the text-to-GUI retrieval performance of our UIClip model is evaluated together with baseline models on three distinct datasets. During the evaluation, every screenshot and its associated captions in test dataset are embedded using the models under evaluation. For every single caption, all screenshots within the test dataset are ranked by the cosine similarity of the caption embedding and screenshot embedding. A comparative analysis of the performance of UIClip against various baselines provide answers to RQ1.

### v2

@@ 2.3.1. Object Detection Model
We use DETR (DEtection TRansformer) (Carion et al., 2020) as screenshot detector. DETR is an end-to-end object detector, composed of a Convolutional Neural Network (CNN) backbone and a encoder-decoder transformer (Vaswani et al., 2017). Atop the transformer decoder, dual output heads are appended: a linear layer dedicated to categorize class labels and a multi-layer perceptron (MLP) tasked with the generation of bounding boxes for the object location. The model uses so-called object queries to detect objects in an image. Each object query is designed to search specifically for one particular object of interest in the given image. We set the number of queries to 1, as we are interested in the largest screenshot within a surrounded screenshot.
2.3.2. Creation of Training Data
We created a separate dataset to fine-tune the detection model for screenshot cropping. From the dataset created in Section 2.2.2, we collected all the surrounded screenshots and labelled the screenshot areas with a bounding box. The annotation was performed with Prodigy (Explosion, 2024). For each image, an author drew one bounding boxes that cover the screenshot. The results were reviewed by another author. The majority of the surrounded screenshots feature only a single screenshot. In instances where a single image contains multiple screenshots, annotation was selectively applied only to the most prominent (largest) screenshot. This resulted in a dataset comprising 1,768 annotated images, each including a delineated bounding box that specifies the screenshot area.
@@ 2.3.3. Detection Performance
Intersection over Union (IoU) is a standard metric for object detection tasks. It evaluates the extent of overlap between the predicted bounding box and the ground truth bounding box. If $IoU=0$, it means that there is no overlap between the boxes, while $IoU=1$ means that the overlap is perfect.
@@
We followed the evaluation protocol of the Common Objects in Context (COCO) (Lin et al., 2014), which is a common object detection benchmark. As stated in the protocol, we quantitatively evaluated the object detection performance by calculating precision and recall metrics at various IoU thresholds. Specifically, we computed the precision values at three IoU levels: 0.50, 0.75, and an average over the range from 0.50 to 0.95. Similarly, recall was determined over the IoU range of 0.50 to 0.95.
We used 80% of the data for training and the remaining 20% for evaluation. The detector was trained on the training set using mini-batch gradient descent, with a batch size of 16 and AdamW optimiser with an initial learning rate of $1e^{-5}$ and ten training epochs. Table 2 shows the results for a 10-fold cross-validation, indicating a top performance.
Table 2. Accuracy of the screen localization (mean values for 10-fold cross-validation).
@@ 2.3.4. Applying the Detector
We trained a DETR model with all of the 1,768 labelled images, and applied this model to the 191,993 images categorized as surrounded screenshot. DETR predicted a bounding box that localize the screenshot for each image. We then cropped these images according to the bounding box. The screenshots cropped from the surrounded screenshot as well as the app introduction images previously classified as screenshots represent the ScreenRepo dataset.
@@ 2.4. Caption Recognition
The majority of the images, classified as surrounded screenshots, include captions that provide a succinct descriptions of the screenshots. Typically, these captions are positioned either above or below the image. In order to accurately extract this textual information, we employed Optical Character Recognition (OCR), a commonly used technology for recognizing text displayed within images.
@@ 2.4.1. Applying OCR
We used PaddleOCR (Li et al., 2022b), which is an industrially robust OCR model renowned for its accuracy and swift performance. The application of PaddleOCR is adequate since the textual content within the images under analysis is markedly legible and usually not hand-written. PaddleOCR processes an input image and outputs a collection of identified elements. Each element is comprised of a bounding box delineating the text area, its recognised text, and an associated confidence level reflecting the recognition certainty. We applied PaddleOCR on all of the 191,993 surrounded screenshots. For each surrounded screenshot, text bounding boxes that exhibit spatial overlap with the screenshot’s bounding box, as delineated in Section 2.3.4, were excluded. The remaining text bounding boxes were subsequently aggregated to form a coherent caption for the corresponding image.
@@ 2.4.2. Post-processing
Some of the surrounded screenshots may not contain captions. These images were removed from our analysis. The captions extracted are not exclusively in English. They encompass a variety of languages including French, German, and Arabic, among others. As we focus on English in this work, captions in languages other than English were excluded. To identify the language of each caption, we employed Lingua (Stahl, 2022), an effective language detection tool. Furthermore, to correct any spelling error within the captions due to potential OCR errors, we utilised Autocorrect (Sondej, 2020), a Python-based spelling corrector.
It is noteworthy that in some instances, multiple surrounded screenshots of an app may feature identical captions. This suggests that the caption is either overly generic or simply a logo of the app, and thus may not adequately convey the screenshot content. To address this redundancy, we conducted a filtering process whereby, for each app, only the first surrounded screenshot from the set containing identical captions was retained. We observed that the first surrounded screenshot displayed by an app on Google Play typically showcases the most representative feature. The post-processing finally left 135,357 screenshot-caption pairs, collectively referred to as the SCapRepo dataset.
@@ 3. GUI Search Engine
We introduce GUing, an advanced search engine based on a vision-language model that leverages textual queries to retrieve relevant screenshots from a large GUI repository which consists of ScreenRepo, Rico (Deka et al., 2017) and ReDraw (Moran et al., 2020). The architecture of GUing is illustrated in Figure 4. GUing is based on the GUIClip model. The engine uses the image encoder and text encoder modules of GUIClip to embed screenshots and textual queries within a unified latent space. This enables an efficient search for the screenshots by calculating the cosine similarity between the text query embedding and the screenshot embedding. In the following, we introduce the details of GUIClip and GUing.
Figure 4. Overview of GUing, our GUI search engine for text-to-GUI retrieval.
@@
Overview of GUing, our GUI search engine for text-to-GUI retrieval.
3.1. Constructing the GUIClip Model
GUIClip, our vision-language model, serves as the fundamental component of the GUI search engine, integrating the image and text modalities. Building on CLIP (Radford et al., 2021), GUIClip offers enhanced capabilities in multimodal representation learning for the GUI domain. CLIP (Contrastive Language–Image Pre-training) is a significant foundation model in multimodal representation learning trained on a large-scale image-caption dataset. In this work, we construct a GUI-specific CLIP model by training a vision-language model on a large-scale dataset of screenshot-caption pairs.
@@ 3.1.1. Training Data and Pre-processing
To train GUIClip learn the visual representation of mobile screenshots, we combined three datasets:
  • Screen2Words dataset (Wang et al., 2021) is a subset of Rico. It contains 112,085 textual summaries for 22,417 unique screenshots. The screenshot summaries are written by professional annotators. For each screenshot, five summaries are created by five different annotators.
  • Clarity dataset (Moran et al., 2022) consists of 45,998 descriptions for 10,204 screenshots from popular Android apps. Crowd workers summarised each screenshot in different granularity, with one high-level caption and up to four low-level detailed descriptions.
  • Google Play Screenshot Caption (SCapRepo) dataset, where textual descriptions are created by developers and app vendors. The dataset includes 135k screenshot-caption pairs (see Section 2). The screenshots are cropped from app introduction images on Google Play. The corresponding captions briefly describe the screenshots.
The captions were first tokenised with the CLIP tokeniser (OpenAI, 2021a). Captions longer than 77 tokens were truncated to 77, captions shorter were padded to 77. To process the screenshot images, we resized the resolution to $224\times 224$. Furthermore, we proportionately rescaled the pixel value of the images from a range of 0-255 to 0-1. This new value was then normalised using the parameters (means and standard deviations) featured in the official CLIP documentation (OpenAI, 2021a).
@@ 3.1.2. Model Architecture
GUIClip consists of two components, an image encoder and a text encoder. The image encoder, which adopts a Vision Transformer (ViT) model (Dosovitskiy et al., 2021), ingests image data and generates its corresponding image embedding. The text encoder is a transformer encoder (Vaswani et al., 2017) which accepts text as input to produce an associated text embedding. Employing contrastive learning techniques, the image and text embeddings are mapped into a multi-modal embedding space. In this shared space, images and texts presenting semantic similarity are mapped close to each other.
Contrastive learning on image-text pairs constructs a bridge between visual perception and natural language. Given a batch of N (image, text) pairs, the training objective is to identify the authentic pairings from $N\times N$ potential (image, text) combinations within the batch (Radford et al., 2021). To achieve this objective, the model formulates a multi-modal embedding space, concurrently training an image encoder and text encoder to maximise the cosine similarity of the image and text embeddings derived from the N authentic pairs in the batch. Simultaneously, it minimises the cosine similarity of the embeddings from the remaining $N^{2}-N$ incorrect pairings.
Figure 5. Overview of contrastive learning to train our vision-language model GUIClip (adapted from (Radford et al., 2021)).
@@
Overview of contrastive learning to train our vision-language model GUIClip (adapted from (Radford et al., 2021)).
@@ 3.1.3. Training Details
In order to adapt the CLIP model to GUIClip, rather than training the CLIP model from scratch where the weights of image encoder and text encoder are randomly initialised, our training is based on the checkpoint of ”openai/clip-vit-base-patch32” (OpenAI, 2021b). This checkpoint has been pre-trained on a 400 million image-text pairs. This method allows us to harness the benefits inherent to pre-trained models and, as a result, substantially diminishes the time required for training. The model was trained on all of the collected screenshot-caption pairs for 5 epochs. We applied mini-batch gradient descent with a batch size of 128 and AdamW optimiser with an initial learning rate of $5e^{-5}$.
@@ 3.2.1. Processing the GUI Repository
The GUI repository is composed of 383k screenshots from ScreenRepo (303k), Rico (66k) and ReDraw (14k). These screenshots are stored in image formats such as .jpeg and cannot be directly queried. Therefore, we embedded all the screenshots in the repository and saved them in a vector database.
The screenshots are first processed (i.e. resized, rescaled, and normalised) with the same parameters discussed in Section 3.1.1. The processed screenshots subsequently serve as input for the image encoder of GUIClip. The image encoder accepts an image as an input and returns a 512-dimensional embedding. The entire GUI repository, comprising 383k screenshots, is processed via the image encoder, consequently generating 383k 512-dimensional embeddings. The mapping between the screenshot ids and their embeddings are saved in a vector database for the retrieval.
@@ 3.2.2. Querying on the GUI Repository
In the querying phase, the textual query is encoded by GUIClip’s text encoder, resulting in a 512-dimensional embedding. Subsequently, this query embedding is utilised to identify the top-k most similar screenshot embeddings from the vector database. We use cosine similarity as the measure of similarity. Consequently, the query results consist of the top-k most similar screenshots.
@@ 3.2.3. Accelerating the Searching
Given the large scale of the GUI repository, employing a brute force approach to compare the query embedding with all screenshot embeddings is inefficient. To tackle this problem, we applied approximate nearest neighbour search (Arya et al., 1998) to accelerate the querying. To this end, we established n Voronoi cells within the embedding space, with each screenshot embedding falling into one of these designated cells. During a search, the query embedding is initially compared with the centroid embeddings of each cell. This method results in a considerable reduction of necessary comparisons when contrasted with the total number of screenshot embeddings. The k cells with centroids closest to the query embedding are identified, whereupon the screenshot embeddings within these cells are thereafter compared with the query embedding in order to pinpoint similar screenshot embeddings. Our implementation is based on Faiss (Douze et al., 2024), a library for efficient similarity search. We set the number of cells n to 3000, and the number of cells that are visited to perform a search k to 1000.
@@ 4. Evaluation Design
For the empirical evaluation of our search engine, we focus on answering the following research questions:
  • RQ1: How does GUing perform in text-to-GUI retrieval tasks compared to state-of-the-art approaches? As the CLIP model (Radford et al., 2021) was trained on 400 millions of image-caption pairs from the internet, it can also be applied on text-to-GUI retrieval task. An alternative method involves leveraging only the textual information: i.e. the text displayed on the GUI images, computing its semantic similarity with the input textual query. Thus, the question is whether the performance our proposed GUIClip model surpasses the CLIP model and text-only approaches.
  • RQ2: How relevant are the retrieved screenshots for queries representing app features? As a search engine, GUing processes textual queries as input and searches for relevant images within the GUI repository. It is thus important to evaluate how effective is GUing in delivering useful results that actually align with user expectations.
To answer these research questions, we designed two experiments: One (Exp1) consists of benchmarking three screenshot-caption datasets, and the other (Exp2) is based on a manual evaluation of the search engines.
4.1. Exp1: Evaluation of Model Performance on Three Datasets
In the first experiment, the text-to-GUI retrieval performance of our GUIClip model is evaluated together with baseline models on three distinct datasets. During the evaluation, every screenshot and its associated captions in test dataset are embedded using the models under evaluation. For every single caption, all screenshots within the test dataset are ranked by the cosine similarity of the caption embedding and screenshot embedding. A comparative analysis of the performance of GUIClip against various baselines provide answers to RQ1.